Wondering if someone could please help me on a Physic problem???
If so here is the question:
A student throws a ball vertically downward at a velocity of 8.0m/s from the top of a 25m building.
a)How long does it take the ball to reach the ground?
b)What is the velocity of the ball as it hits the ground?
Please & Thank you if someone could help me???

Answer by mananth(16946)

A student throws a ball vertically downward at a velocity of 8.0m/s from the top of a 25m building.
a)How long does it take the ball to reach the ground?
b)What is the velocity of the ball as it hits the ground?
s=V0t +1/2 gt^2 (g=9.8 m/s)
25=8t+4.9t^2
4.9t^2+8t-25=0
use quadratic formula
you will get t = 1.585 s or 1.6 seconds
...
v=u+at
v=8+9.8*1.6
23.68m/sec